OpenStudy (anonymous):

Find the cross product and show it is orthogonal to both u and v. u= 6k v= -i+3j+k

OpenStudy (anonymous):

This is the answer I got. I dont know if its correct or not. Can someone help?

OpenStudy (amoodarya):

you are in right way but for orthogonality you may show u*v =||u|| ||v|| sin (u,v) find sin (u,v) it must be 1 because sin pi/2=1

OpenStudy (anonymous):

Can you show me how? I want to make sure I am doing the steps correctly.Also, is the answer I got correct?

OpenStudy (amoodarya):

 name u*v=a find a.u it will be zero then you shoe a has right angle with u also do for a and v

OpenStudy (anonymous):

So the answer is..... sqrt -10j -------- 10

OpenStudy (amoodarya):

u*v=-18 i -6j +0k (u*v).u=(-18 i -6j +0k).( 6k )=0 so (u*v).v=(-18 i -6j +0k) .(-i+3j+k )=+18-18 =0

OpenStudy (amoodarya):



OpenStudy (anonymous):

I dont think that is the cross product? That is what it is asking me to find. Im confused :0

OpenStudy (amoodarya):

you calculate the cross product correct u*v=-18i -6j +0k so now to show orthogonality between u*v and v use the dot product if dot product goes to zero so they are orthogonal

OpenStudy (anonymous):

The answer is 1?

OpenStudy (anonymous):

I got that the cross product is 0.